OpenStudy (anonymous):

19=h/3rds -8 what does h=

OpenStudy (compassionate):

\[19 = \frac{ 3 }{ -8 }\] Is this the right equation?

OpenStudy (anonymous):

no it is 19=-8

OpenStudy (anonymous):

if you got that

OpenStudy (anonymous):

btw thank you for helping

OpenStudy (compassionate):

h = 81

OpenStudy (anonymous):

i think it would be h=81rds

OpenStudy (anonymous):

how do you do the problem ??? please xexplain

OpenStudy (anonymous):

explain*

OpenStudy (compassionate):

\[19 = \frac{ h }{ 3 } - 8\] Add eight to both sides: \[27 =\frac{ h }{ 3 }\] Multiply by 3. \[81 = h \]

OpenStudy (anonymous):

19 = h/3rds - 8 +8 +8 27 = h/3rds x3rds x3rds 81rds = h

OpenStudy (compassionate):

As always, medals are appreciated.

OpenStudy (anonymous):

thabk you both

OpenStudy (anonymous):

compassionate is doing it right but shes not putting the rds

OpenStudy (anonymous):

rds was just to tell you thirds sorry